# MultiBanana: A Challenging Benchmark for Multi-Reference Text-to-Image Generation

### v1

@@ 3.2 How to Construct MultiBanana
Table 3: Average performance per model across different task categories (10-point scale; the higher the better). The average scores from Gemini-2.5 and GPT-5 are reported. Both open-source and closed-source models exhibit notably lower performance on background replacement tasks, regardless of the number of reference images.
@@
3.3 Statistics in MultiBanana
In the real dataset, the largest major category is background, comprising 798 images, one-fourth of the total. These mainly consist of images without distinct subjects. The next largest categories are person (741 images) and style (579 images), and together these three categories account for 70% of the dataset. In contrast, light (195 images), object (413 images), and tone (118 images) are relatively small in size. Although these categories are useful as references for changing global conditions such as illumination and overall atmosphere, their sample sizes remain limited.
Within the object category, there are 111 images of animals, 40 of clothing and textile design, 71 of industrial design, and 191 of miscellaneous objects, indicating that data focusing on specific materials or design patterns is scarce. Most images in the background, style, and tone categories lack clear subjects. Furthermore, images containing textual elements, such as logos or signs, are nearly absent. The person category is subdivided into finer attributes: face expression (10), female (181), male (139), hairstyle (135), makeup (59), and pose usefulness (217). While pose-related data account for the largest portion (around 30%), reference data capturing subtle characteristics such as facial expression or makeup remain limited, resulting in a skewed distribution.
To address this imbalance, synthetic data were generated, focusing primarily on images containing explicit subjects (e.g., persons, animals, and objects) as well as those including text. The generation process was designed to intentionally expand the coverage of underrepresented subcategories and improve the overall balance among attributes.
[…]
Figure 5: Changes in scores for each evaluation criterion when varying the number of reference images. Both open-source and closed-source models exhibit a general trend of decreasing all scores as the number of references increases.
Task Statistics.   After filtering the generated editing instructions using both Gemini and manual checks, we obtained 3,769 high-quality reference images and instruction sets. Among them, 264 sets correspond to single-reference tasks, 907 to two-reference tasks, and 2,598 to multi-reference tasks. Figure 4 provides a further breakdown of the multi-reference tasks, showing that they are evenly distributed with respect to the number of reference images. Additionally, for each task type, X Object, X–1 Object + Background, X–1 Object + Local, and X–1 Object + Global, we ensured that at least 70 sets were included. This is substantially larger than the number of sets in multi-reference tasks of DreamOmni2, which is on the order of ten, enabling more reliable evaluation of model performance. Furthermore, Table 2 shows that our newly introduced evaluation axes – cross-domain, scale and viewpoint differences, rare concept, and multilingual – provide sufficient samples. See Appendix B for details.
Figure 6: Results for diffucult reference combinations. For cross-domain and different scale and view tasks, every model shows lower scores than tasks without such conditions.
@@ 3.4 Evaluation Setting
For evaluating the quality of edited images, human evaluation is among the most valuable methods, yet gathering human feedback at scale is prohibitively costly. A practical approach to reduce time and cost is to leverage AI feedback from VLMs (e.g., Gemini-2.5 [geminiteam2023gemini] and GPT-5 [openai2025gpt5]), which has been shown to modestly align with human judgments of image quality [na2024boost, wu2024boosting, furuta2024improving, oshima2025inference]. Therefore, we used Gemini-2.5 and GPT-5 to rate all generated images.
@@ 4 Experiments
We adopt several closed models (i.e., Nano Banana [google2025nanobanana], GPT-Image-1 [openai2025gpt4oimage]) and open models (i.e., Qwen-Image-Edit-2509 [wu2025qwenimagetechnicalreport]], DreamOmni2 [xia2025dreamomni2], OmniGen2 [wu2025omnigen2]) for MultiBanana evaluation. For all models except OmniGen2, we generate outputs for every combination of reference images and text prompts included in the MultiBanana benchmark. Since OmniGen2 does not support more than six reference images, we only evaluate it under the 5-reference setting. As described in Section 3.4, the evaluation is conducted using Gemini-2.5 [geminiteam2023gemini] and GPT-5 [openai2025gpt5]. In the main paper, we report the average scores from Gemini-2.5 and GPT-5, while the individual results for each evaluator are provided in the supplementary material.
Figure 7: (Left) Failure cases of cross-domain, different scale, rare concept, multilingual text. (Right) Failure cases of multi-references in open-source models. They tend to ignore multiple subjects under high-reference conditions.
@@ 4.1 Per-Task Evaluation
We compute the MultiBanana scores for each task type defined in Section 3.1, and present the results in Table 3 and Section E.1. For the two-reference tasks, we report the average score across all 11 task variants. For the multi-reference tasks, we compute the average score for each task type (i.e., X Objects, X–1 Objects + Local, X–1 Objects + Global, X–1 Objects + Background) across $X=\{3,4,5,6,7,8\}$.
Overall, both open-source and closed-source models exhibit notably lower performance on background replacement tasks, regardless of the number of reference images. The results for the two-reference setting also reveal clear performance gaps between closed-source and open-source models, even for relatively simple editing scenarios such as single-reference or two-reference tasks.
@@ 4.2 Effect of the Number of References
We investigate how the MultiBanana score changes when varying the number of reference images $X=\{3,4,5,6,7,8\}$ for each multi-reference generation task (Figure 5). Both open-source and closed-source models exhibit a general trend of decreasing Total Score as the number of references increases. However, the nature of this degradation differs markedly between the two model families.
Closed-source models demonstrate strong adherence to reference attributes and prompt instructions. As a result, even when the number of references grows, the decline in Instruction Alignment and Reference Consistency remains relatively mild. Nevertheless, because these models attempt to satisfy all reference constraints strictly, increasing the number of subjects often leads to degraded overall visual quality, including over-crowded scenes and compositional inconsistencies (see Figure 1 and Figure 7; left).
In contrast, open-source models show weaker adherence to prompts and references; even under the 3-reference setting, both Instruction Alignment and Reference Consistency remain low, and these scores further deteriorate as the reference count increases. Qualitative inspection also reveals that open models frequently ignore multiple subjects under high-reference conditions. However, their overall quality (background consistency, physical realism, visual quality) remains relatively stable, and they tend not to suffer from the severe compositional collapse observed in closed models (see Figure 7; right).
In summary, closed-source models succeed in incorporating all required subjects but often produce globally inconsistent or distorted scenes due to overfitting to fine-grained reference details. Open-source models, by contrast, tend to omit multiple reference subjects in complex scenarios but generate visually clean images with fewer structural failures. These observations highlight a clear trade-off between strict reference fidelity and holistic image coherence.
@@ 4.3 Analysis of Difficult Reference Combinations
Next, to analyze the challenging subject combinations included in this benchmark, namely those described in Section 3.1, we compare the MultiBanana scores between tasks containing these difficult combinations and those that do not in Figure 6. For cross-domain, different-scale and view, rare-concept, and multilingual tasks, most models show lower scores than for tasks without such conditions. These results indicate that the difficult reference concepts focused on by our benchmark remain challenging, even for state-of-the-art closed models. This demonstrates that our benchmark effectively uncovers the current limitations of multi-reference image generation.
@@ 4.4 Agentic Framework
To address the performance limitations on our challenging benchmark, we further construct three agentic framework baselines: Iterative Prompt Refinement (IPR), Context-Aware Feedback Generation (CAFG), and Selective Reference Adaptation (SRA). We then conduct experiments on our MultiBanana benchmark. We show the results of IPR in Table 3, where the generator creates images, and based on these images, the planner rewrites the prompt over multiple steps. While GPT’s performance improved with the agent framework, Nano Banana showed no improvement and, in some cases, performance degraded due to information loss from the initial instruction. For details on the problem setting, ablations, and results, see the supplementary material.
@@ 5 Conclusion
We constructed a comprehensive benchmark for multi-reference image generation, focusing on challenges specific to this setting, such as difficult reference combinations. Using it, we evaluated state-of-the-art open and closed models and found two notable failure modes as the number of references increases: models either ignore editing instructions or collapse when trying to follow them. Across different types of reference difficulty, tasks involving cross-domain inputs, large-scale or viewpoint gaps, rare concepts, and multilingual inputs are consistently challenging.
@@ Acknowledgements
We thank Minsu Kim and Heiga Zen for their support on this work and review on the initial version of the paper. We also appriciate the funding support from Google Japan. MS was supported by JSPS KAKENHI Grant Number JP23H04974.
[…]
@@
A Details on Benchmark Construction
A.1 Details on Image Collection
@@
We constructed the reference image collection for our benchmark using both real images sourced from LAION-5B [laion2022aesthentic] and synthetic images generated by Nano Banana [google2025nanobanana] and GPT-Image-1 [openai2025gpt4oimage]. By combining real and synthetic data, we mitigated the categorical biases inherent in the real-image distribution while substantially expanding the coverage and diversity of the reference set. To generate synthetic images, we designed distinct axes of variation for four major categories: humans, animals, objects, and text-containing images.
@@
Human Category  We systematically combined attributes such as makeup, hairstyle, hair color, pose, strong facial expressions, lighting, and overall tonal style to create a wide range of appearance conditions. Examples include ethnic or fashion-model-style makeup, vivid or unconventional hair colors, model-like or manga-like poses, strong emotional expressions (e.g., anger, joy, surprise), dramatic lighting such as window light through blinds or neon illumination, and global color tones such as sepia or bluish palettes. These conditions were intentionally selected to supplement variations that are underrepresented in real-world data.
@@
Animal Category  We first specified major species groups (dogs, cats, birds, reptiles, wild animals, and fantasy animals). Then we added attributes to enhance the diversity specific to animal imagery. These included human-like behaviors (e.g., a bear reading a book, a fox playing in a rock band), rare colors (e.g., a blue horse or a pink panda), distinctive lighting conditions, and varied global tones such as sepia. This incorporated rare or extreme examples that seldom appear in real datasets.
@@
Object Category  We expanded diversity along two axes: intrinsic object attributes and photographic conditions. The former included distinctive textures (e.g., denim, leather, knit), characteristic prints (e.g., floral, striped, checkered), and variations in material or design (e.g., matte, glossy, transparent). The latter included variations in lighting and global tonal adjustments (e.g., sepia, cool color tones), thereby ensuring a broad range of appearance conditions for objects.
@@
Images Containing Text  Images containing text, which were largely absent from real data, were generated exclusively using GPT-Image-1, as Nano Banana exhibited limited ability to render non-English text. We defined three levels of text length (a single word, a short phrase of up to five words, and a longer sentence). We combined each with two scene types: text rendered directly within a landscape (e.g., ocean, forest) and text appearing naturally in the environment (e.g., on signs in real scenes). These six prompt patterns were generated in English and subsequently translated into Japanese and Chinese to enrich multilingual text-containing images.
@@
Scene Diversity  Finally, all synthetic prompts were designed to incorporate scene diversity, including European cityscapes, university campuses, tropical beaches, event halls, and other varied environments. The generated subjects were ensured to be well-framed, resulting in images suitable for use as high-quality reference material.
A.2 Details on Image Filtering
First, to remove images unsuitable for editing, the images were filtered based on the size of the foreground objects, following the manner in [ye2025imgedit]. YOLOv12 [tian2025yolov12] was applied to all images to detect objects and then performed semantic segmentation using SAM [ravi2025sam], conditioned on the detected bounding boxes. Images were discarded if they met any of the following criteria: the bounding box area covered less than 2% of the entire image, the segmented region within the bounding box covered less than 30%, or the CLIP similarity [ma2022xclip] between the YOLO-predicted class name and the bounding box region was below 20. These cases were judged to contain objects that were not clearly visible. However, images where no objects were detected were retained, assuming they are useful for background or style-level editing tasks. After this filtering, 48% of the images were retained.
We also filtered out images that were inappropriate or unsuitable as references, such as unsafe content, charts, and screenshots of system messages, using both Gemini-based automated screening and human review. For inappropriate content, we specifically targeted categories including hate, harassment, violence, self-harm, sexual content, nudity, shocking content, illegal activity, and other distressing material, ultimately excluding approximately 3% of the collected images. We also manually removed synthetic data that was nearly identical. An example is shown in Figure 8.

### v2

@@ 3.3 Statistics in MultiBanana
Table 3: Average performance per model across different task categories (10-point scale; the higher the better). The average scores from Gemini 2.5 and GPT-5 are reported. Both open-source and closed-source models exhibit notably lower performance on background replacement tasks, regardless of the number of reference images.
@@
Figure 5: Changes in scores for each evaluation criterion when varying the number of reference images. Both open-source and closed-source models exhibit a general trend of decreasing scores across all metrics as the number of references increases.
@@ 3.4 Evaluation Setting
For evaluating the quality of edited images, human evaluation is among the most valuable methods, yet gathering human feedback at scale is prohibitively costly. A practical approach to reducing time and cost is to leverage AI feedback from VLMs (e.g., Gemini 2.5 [19] and GPT-5 [57]), which has been shown to align reasonably with human judgments of image quality [38, 53, 82, 17, 58]. Therefore, we used Gemini 2.5 and GPT-5 to rate all generated images.
@@ 4 Experiments
We adopt several closed models (i.e., Nano Banana [22], GPT-Image-1 [56]) and open models (i.e., Qwen-Image-Edit-2509 [79], DreamOmni2 [83], OmniGen2 [80]) for MultiBanana evaluation. For all models except OmniGen2, we generate outputs for every combination of reference images and text prompts included in the MultiBanana benchmark. Since OmniGen2 does not support more than 6 reference images, we evaluate it only under the 5-reference setting. As described in Section 3.4, Gemini 2.5 [19] and GPT-5 [57] are used for evaluation. In the main paper, we report the average scores from Gemini 2.5 and GPT-5. See Appendix A for the results of Qwen3-VL as judge, and Appendix F for the detailed results and discussion.
Figure 6: (Left) Failure cases of closed-source models under challenging conditions such as cross-domain references, scale mismatches, rare concepts, and multilingual text. (Right) Failure cases of open-source models under many-reference conditions. They tend to ignore some reference subjects entirely.
@@ 4.1 Per-Task Evaluation
We compute the MultiBanana scores for each task type defined in Section 3.1, and present the results in Table 3. For the two-reference tasks, we report the average score across all 11 task variants. For the multi-reference tasks, we compute the average score for each task type (i.e., X Objects, X–1 Objects + Local, X–1 Objects + Global, X–1 Objects + Background) across $X=\{3,4,5,6,7,8\}$.
Both open-source and closed-source models exhibit lower performance on background-replacement tasks, regardless of the number of reference images. The results also reveal clear performance gaps between closed-source and open-source models, even in relatively simple editing scenarios, such as single- and two-reference tasks.
@@ 4.2 Effect of the Number of References
We investigate how the MultiBanana score changes when varying the number of reference images $X=\{3,4,5,6,7,8\}$ for each multi-reference generation task (Figure 5). Both open-source and closed-source models exhibit a general trend of decreasing Total Score with increasing number of references. However, the nature of this degradation differs between the two model families.
Closed-source models demonstrate strong adherence to reference attributes and prompt instructions. As a result, even when the number of references grows, the decline in Instruction Alignment and Reference Consistency remains relatively mild. Nevertheless, because these models attempt to strictly satisfy all reference constraints, increasing the number of subjects often degrades overall visual quality, including overcrowded scenes and compositional inconsistencies (see Figure 1 and Figure 6; left).
In contrast, open-source models exhibit weaker adherence to prompts and references; even under the 3-reference setting, both Instruction Alignment and Reference Consistency remain low, and these scores deteriorate further as the number of references increases. Qualitative inspection also reveals that open models frequently ignore multiple subjects under high-reference conditions. However, their overall quality (background consistency, physical realism, visual quality) remains relatively stable, and they tend not to suffer from the severe compositional collapse observed in closed models (see Figure 6; right).
In summary, closed-source models succeed in incorporating all required subjects but often generate globally inconsistent or distorted scenes due to overfitting to fine-grained reference details. Open-source models, by contrast, tend to omit multiple reference subjects in complex scenarios but generate visually clean images with fewer structural failures. These observations highlight a clear trade-off between strict reference fidelity and holistic image coherence.
Figure 7: Results for difficult reference combinations. For cross-domain and different scale and view tasks, every model shows lower scores than tasks without such conditions.
@@ 4.3 Analysis of Difficult Reference Combinations
To analyze the challenging reference combinations described in Section 3.1, we compare MultiBanana scores between tasks containing these difficult combinations and those that do not, as shown in Figure 7. For cross-domain, different scale and view, rare-concept, and multilingual tasks, most models achieve lower scores than on tasks without such conditions. These results indicate that the difficult reference concepts focused on by our benchmark remain challenging, even for state-of-the-art closed models. This demonstrates that our benchmark effectively uncovers the current limitations of multi-reference image generation.
4.4 Reliability of VLM Judge
To assess the reliability of VLM-based evaluation, we measured the correlation between VLM and human ratings using three human evaluators on 32 samples from our benchmark (Table 4). Both GPT-5 and Gemini 2.5 show strong correlation with human judgments. This validates the use of VLM judges as a time- and cost-efficient proxy for human evaluation, as described in Section 3.4. We also confirmed that Qwen3-VL-8B-Instruct [2] is a reliable open-source alternative, thereby broadening accessibility for researchers without API access. See Appendix E for further discussion.
@@ 4.5 Agentic Framework
To address the performance limitations on our challenging benchmark, we further construct three agentic framework baselines: Iterative Prompt Refinement (IPR), Context-Aware Feedback Generation (CAFG), and Selective Reference Adaptation (SRA). We then conduct experiments on our MultiBanana benchmark. We show the results of IPR in Table 3, where the generator creates images, and the planner rewrites the prompt over multiple steps based on these images. While GPT’s performance improved with the agent framework, Nano Banana showed no improvement and, in some cases, performance degraded due to information loss from the initial instruction. For details on the problem setting, ablations, and results, see Appendix F.7.
Table 4: Correlation between human and VLM judges.
Judge Pearson $r$ Spearman $r$ Cohen’s $\kappa$
GPT-5 0.6936 0.7134 0.6135
Gemini 2.5 0.5726 0.6113 0.3908
Qwen3-VL 0.5457 0.5642 0.5235
@@ 5 Conclusion
We constructed a comprehensive benchmark for multi-reference image generation, focusing on challenges specific to this setting, such as difficult reference combinations. Using it, we evaluated state-of-the-art open and closed models and found two notable failure modes as the number of references increased: models either ignored editing instructions or collapsed when trying to follow them. Across different types of reference difficulty, tasks involving cross-domain inputs, large-scale or viewpoint gaps, rare concepts, and multilingual inputs are consistently challenging.
@@ Acknowledgements
We thank Minsu Kim and Heiga Zen for their support with this work and for their review of the initial version of the paper. We also appreciate the funding support from Google Japan. MS was supported by JSPS KAKENHI Grant Number JP23H04974.
@@
A Results of Qwen3-VL Judge
Table 5: Per-task total scores for each single- and two-reference task, evaluated by Qwen3-VL. “Back.” denotes the Background task. Both open-source and closed-source models exhibit lower performance on background replacement tasks. Note that for Nano Banana and GPT-Image-1, we use the versions available as of January 2026. Model Single Add Back. Color Hair Makeup Material Pose Replace Style Text Tone DreamOmni2 8.47 6.51 4.59 8.60 7.40 7.89 6.76 6.25 6.86 5.10 5.48 5.14 OmniGen2 9.04 5.34 6.04 5.07 6.70 6.68 6.12 5.91 6.34 3.24 4.29 4.00 Qwen-Image-Edit-2509 9.16 7.01 4.34 6.46 8.00 6.68 5.86 5.87 7.57 3.41 4.95 6.08 Nano Banana 9.45 8.13 7.97 9.67 9.16 9.25 8.95 8.76 8.30 9.03 8.81 8.52 GPT-image-1 9.51 8.09 8.08 9.70 9.04 9.24 9.17 8.80 8.30 9.10 8.52 8.78
Table 6: Per-task total scores for each multi-reference task, evaluated by Qwen3-VL. The local, global, back, and object columns under X correspond to X–1 Objects + Local, X–1 Objects + Global, X–1 Objects + Background, and X Object, respectively. Both open-source and closed-source models exhibit a general trend of decreasing scores across all tasks as the number of references increases. They also tend to perform worse on background replacement tasks, especially as the number of reference images increases. Note that for Nano Banana and GPT-Image-1, we use the versions available as of January 2026. Model 3-references 4-references 5-references local global back object local global back object local global back object DreamOmni2 4.39 4.27 4.33 3.68 3.80 2.95 2.77 2.90 3.23 2.12 2.86 2.26 OmniGen2 4.71 4.86 4.39 4.13 4.37 3.86 3.43 2.95 3.99 2.75 2.14 2.26 Qwen-Image-Edit-2509 5.22 5.81 5.81 5.63 3.24 4.51 4.42 3.94 1.05 1.48 1.28 1.10 Nano Banana 7.63 7.21 6.96 6.91 7.96 6.77 6.47 6.78 7.96 6.54 6.03 5.87 GPT-image-1 7.83 7.26 6.99 6.72 7.73 6.74 5.94 6.83 7.92 6.54 6.16 5.83 Model 6-references 7-references 8-references local global back object local global back object local global back object DreamOmni2 3.34 2.26 1.72 1.93 2.61 2.45 2.10 1.81 2.44 1.65 1.83 1.59 OmniGen2 - - - - - - - - - - - - Qwen-Image-Edit-2509 1.00 1.49 1.00 1.00 1.00 1.46 1.00 1.13 1.00 1.00 1.00 1.00 Nano Banana 6.84 6.08 5.34 5.64 6.37 5.78 5.13 5.44 6.46 6.24 5.31 5.28 GPT-image-1 7.11 5.86 5.51 5.92 6.81 5.78 5.37 5.74 7.07 5.86 5.14 5.55
In Section 4, we report the average scores from Gemini 2.5 and GPT-5 as judges. As demonstrated in Section 4.4, Qwen3-VL-8B-Instruct [2] achieves strong correlation with human judgments comparable to GPT-5 [57] and Gemini 2.5 [19], confirming its reliability as a judge model (see Table 4). Although we used stable API versions for the closed-source judges, to ensure full reproducibility, we additionally employ Qwen3-VL-8B-Instruct, an open-weight, fixed-version model, as the judge. This also broadens access to our benchmark for researchers who do not have access to closed-source VLMs via APIs.
Table 5 and Table 6 present per-task total scores evaluated using Qwen3-VL as the sole judge, calculated in the same way as described in Section 3.4. The results are broadly consistent with those obtained using closed-model judges. The overall ranking of image generation models is preserved. We also observe the same key trends: performance degrades as the number of reference images increases, and background replacement tasks tend to yield lower scores across both open-source and closed-source models. These observations reinforce the validity of Qwen3-VL as a reliable, reproducible alternative. We expect this fixed, open-weight judge to serve as a dependable baseline for future evaluations on our benchmark.
B Details on Benchmark Construction
B.1 Image Collection
@@
We constructed the reference image collection for our benchmark using both real images sourced from LAION-5B [68] and synthetic images generated by Nano Banana [22] and GPT-Image-1 [56]. By combining real and synthetic data, we mitigated the categorical biases inherent in the real-image distribution while expanding the coverage and diversity of the reference set. To generate synthetic images, we designed distinct axes of variation for four major categories: humans, animals, objects, and text-containing images.
@@
Human Category.  We systematically combined attributes such as makeup, hairstyle, hair color, pose, strong facial expressions, lighting, and overall tonal style to create a wide range of appearance conditions. Examples include ethnic or fashion-model-style makeup, vivid or unconventional hair colors, model-like or manga-like poses, strong emotional expressions (e.g., anger, joy, surprise), dramatic lighting (e.g., window light through blinds or neon illumination), and global color palettes (e.g., sepia or bluish). These conditions were intentionally selected to supplement the variations underrepresented in real-world data.
@@
Animal Category.  We first specified major species groups (dogs, cats, birds, reptiles, wild animals, and fantasy animals). We then added attributes to enhance the diversity of animal imagery. These included human-like behaviors (e.g., a bear reading a book, a fox playing in a rock band), rare colors (e.g., a blue horse or a pink panda), distinctive lighting conditions, and varied global tones such as sepia. This incorporated rare or extreme examples that seldom appear in real datasets.
@@
Object Category.  We expanded diversity along two axes: intrinsic object attributes and photographic conditions. The former included distinctive textures (e.g., denim, leather, knit), characteristic prints (e.g., floral, striped, checkered), and variations in material or design (e.g., matte, glossy, transparent). The latter included variations in lighting and global tonal adjustments (e.g., sepia, cool color tones), thereby ensuring a broad range of appearance conditions for objects.
@@
Images Containing Text.  Images containing text, which were largely absent from real data, were generated exclusively using GPT-Image-1, as Nano Banana exhibited limited ability to render non-English text. We defined three levels of text length (a single word, a short phrase of up to five words, and a longer sentence). We combined each with two scene types: text rendered directly within a landscape (e.g., ocean, forest) and text appearing naturally in the environment (e.g., on signs in real scenes). These six prompt patterns were generated in English and subsequently translated into Japanese and Chinese to enrich the multilingual text in images.
@@
Scene Diversity.  Finally, all synthetic prompts were designed to incorporate scene diversity, including European cityscapes, university campuses, tropical beaches, event halls, and other varied environments. The generated subjects were well-framed, resulting in images suitable for use as high-quality reference material.
@@
B.2 Image Filtering
First, to remove images unsuitable for editing, we filtered them based on the size of the foreground objects, as in [91]. YOLOv12 [73] was applied to all images to detect objects and then performed semantic segmentation using SAM [64], conditioned on the detected bounding boxes. Images were discarded if they met any of the following criteria: the bounding box area covered less than 2% of the entire image, the segmented region within the bounding box covered less than 30%, or the CLIP similarity [48] between the YOLO-predicted class name and the bounding box region was below 20. These cases were judged to contain objects that were not clearly visible. However, images where no objects were detected were retained, assuming they are useful for background or style-level editing tasks. After this filtering, 48% of the images were retained.
We also filtered out images that were inappropriate or unsuitable as references, such as unsafe content, charts, and screenshots of system messages, using both automated screening based on Gemini and human review. For inappropriate content, we specifically targeted categories including hate, harassment, violence, self-harm, sexual content, nudity, shocking content, illegal activity, and other distressing material, ultimately excluding approximately 3% of the collected images. We also manually removed nearly identical synthetic data. An example is shown in Figure 8.